OpenStudy (anonymous):

Write the equation of a line with slope m = 6 and including point (0, 1)

OpenStudy (anonymous):

Slope intercept form is y=mx+b right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Put 6 in for m and 1 in for y. You also need to put 0 in for x. So now you have the equation 1=6(0)=b. Can you solve it now?

OpenStudy (anonymous):

Oops, I mean 1=6(0)+b

OpenStudy (anonymous):

You should get a slope of -1.

OpenStudy (anonymous):

Sorry, i'm horrible at math. I tried writing it down on a piece of paper and am still so confused. I have 4 answers to choose from.

OpenStudy (anonymous):

What are your answers?

OpenStudy (anonymous):

A. y = 6x - 1/2 B. y=6x +1/2 C. y = 6x + 1 D. y = 6x - 1

OpenStudy (anonymous):

y=6x-1 is your answer.

OpenStudy (anonymous):

Thank you so much!